Maximum Bank has analyzed the accounts receivable of Scientific​ Software, Inc. The bank has chosen eight accounts totaling ​$210,000 that it will accept as collateral. The​ bank's terms include a lending rate set at prime +3​% and a 2​% commission charge. The prime rate currently is 8​%.

a.  The bank will adjust the accounts by 10​% for returns and allowances. It then will lend up to 95​% of the adjusted acceptable collateral. What is the maximum amount that the bank will lend to Scientific​ Software?

b.  What is Scientific​ Software's effective annual rate of interest if it borrows $135,000 for 12 months? For 6 ​months? For 3 months?  ​(​Note: Assume a​ 365-day year and a prime rate that remains at 8​% during the life of the​ loan.)
Answer #1

ma Given that 0.40 Bank chosen cight accounts totaling - $210,000 Habe set at Purime +3y. fam. cuvent porime fabe. = 8y. (9)-) 17550 135,000 =0:139ers 13% (2) Effective annual rate of interest: Co moniche = (tote Interest expense = (133,000 X 2y.)+(Effective annual rate a = (1.04754-1 = 1.2039 - 1 +0.2039 =120.39% Conclusion → Effective annual rate of Interest (amount bor